OpenStudy (anonymous):

write the polynomial in factored form, 4x^3+8x^2-96x @Little_Kay

OpenStudy (anonymous):

=4x(x2+2x-24) =4x(x2+6x-4x-24) =4x{x(x+6)-4(x+6)} =4x(x+6)(x-4)

OpenStudy (anonymous):

what are the zeroes of the function? graph the function. y=x(x+2)(x+5)

OpenStudy (anonymous):

set the the equations in parenthesis equal to 0. x+2=0, x+5=0. can you solve that?

OpenStudy (anonymous):

yes i got the answer to number 2 thanks do you want me to start in 3 in a different chat thing?

OpenStudy (anonymous):

Yeah hold on though. what did you get for your zeros ?

OpenStudy (anonymous):

is it 2 and 5?

OpenStudy (anonymous):

probably not im horrible at doing those

OpenStudy (anonymous):

Nope. when you set x+2=0 you have to subtract 2 from both sides. so 0-2=-2, 0-5=-5 so your zeros would be 0,-2,-5

OpenStudy (anonymous):

oh i see, i didn't know i hald to do that to them! lo

OpenStudy (anonymous):

Lol go ahead and close this one